Stanley and Phillys leave the office and travel in opposite directions. Stanley drives 15 mph slower than Phyllis, and after 2 hr they are 146 mi apart. How fast was each person driving?
Answer by mananth(16949)

Phyllis x mph
Stanley x+ -15
..
They are going away from each other.

So add up their speed.

combined speed = x+x-15
(2x-15 )
Time = 2 hours
Distance = 146 miles
Distance = speed * time
( 2 x -15 )* 2 = 146
4 x -30 = 146
4 x = 146 + 30
4 x = 176
/ 4
x= 44 mph Phyllis
Stanley 44 + -15 = 29 mph